How does the equation for valuing a $1000 par value bond change if semiannual payments are made? Find the value of a 10-year, semiannual payment, 10 percent coupon bond if nominal rd = 13%.
Answer #1

If semi annual payments are made:

Number of periods = 10 * 2 = 20

Coupon = 0.1 * 1000 = 100 / 2 = 50

Rate = 13% / 2 = 6.5%

Value of Bond = Annuity * [ 1 - 1 / ( 1 + r)n] / r + FV / ( 1 + r)n

Value of Bond = 50 * [ 1 - 1 / ( 1 + 0.065)20] / 0.065 + 1000 / ( 1 + 0.065)20

Value of Bond = 50 * 11.018507 + 283.797029

Value of Bond = $834.72